OYmason:

Use the rule x+9 to write a sequence of numbers. Then write the sequence as a list of ordered pairs. Start with x=0 and substitute at least six values for

Aratox:

Start with x=0 and substitute at least six values for ??

umm:

Welcome to QuestionCove! Please post the full question for us to be able to answer your question correctly, thanks!

OYmason:

I'll just screenshot

Aratox:

Alright, that gives us enough information to reply, I'll start working on it right away. Using the rule x+9, let's substitute at least six values for x and write the corresponding sequence of numbers as a list of ordered pairs. 1) When x = 0, the number in the sequence is: 0 + 9 = 9. The ordered pair is (0, 9). 2) When x = 1, the number in the sequence is: 1 + 9 = 10. The ordered pair is (1, 10). 3) When x = 2, the number in the sequence is: 2 + 9 = 11. The ordered pair is (2, 11). 4) When x = 3, the number in the sequence is: 3 + 9 = 12. The ordered pair is (3, 12). 5) When x = 4, the number in the sequence is: 4 + 9 = 13. The ordered pair is (4, 13). 6) When x = 5, the number in the sequence is: 5 + 9 = 14. The ordered pair is (5, 14). 7) When x = 6, the number in the sequence is: 6 + 9 = 15. The ordered pair is (6, 15). So, the sequence of numbers as a list of ordered pairs is: (0, 9), (1, 10), (2, 11), (3, 12), (4, 13), (5, 14), (6, 15). If you have any more questions, please feel free to ask me!

toga:

Using the rule x+9, we can write the following sequence of numbers starting with x=0: 0, 9, 18, 27, 36, 45 To write the sequence as a list of ordered pairs, we can pair each term with its corresponding value of x. Therefore, the ordered pairs would be: (0,0), (1,9), (2,18), (3,27), (4,36), (5,45) I hope this helps! Let me know if you have any other questions.

pup:

9

shelovejayy:

9